OpenStudy (anonymous):

(1 pt) What property is shown by the equation? Column A Column B 1. –(–33 + 18) = 33 + (–18) 2. –10 + (–15) = –15 + (–10) 3. 4 + (29 + 16) = 4 + (16 + 29) 4. –12 + (12 + 27) = (–12 + 12) + 27 A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

@Ms-Brains @Mehek14

OpenStudy (anonymous):

@Mehek14 @Ms-Brains

OpenStudy (anonymous):

@Sushi121212

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Alrighty, I can help you. :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In this case, it would be associative. \ Here's why: Although, they can be arranged in many ways and gave the same answer. It is NOT commutative, because they are undergoing changes in value and re-arrangement. Did that help? :)

OpenStudy (anonymous):

a,b,c 1.–(–33 + 18) = 33 + (–18) a,b,c 2.–10 + (–15) = –15 + (–10) a,b,c 3.4 + (29 + 16) = 4 + (16 + 29) a,b,c 4.–12 + (12 + 27) = (–12 + A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

for each you have to chose a b or c @k-nowhere

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Hmm, I know answers 1-3, but it seems part of #4 has been cut off? "-12+

OpenStudy (anonymous):

oh 1 sec

OpenStudy (anonymous):

4. –12 + (12 + 27) = (–12 + 12) + 27

OpenStudy (anonymous):

?

OpenStudy (anonymous):

1.) Would be B, the explanation is the same is the one we did before. 2.) Would be A, because the values themselves aren't changing; just the order. 3.) Would be A, because only the order of the values or changing 4.) Would be A, same ^

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem :)